# Jamming is a first-order transition with quenched disorder in amorphous materials sheared by cyclic quasistatic deformations

### v2

@@ Abstract
Jamming is an athermal transition between flowing and rigid states in amorphous systems such as granular matter, colloidal suspensions, complex fluids and cells. The jamming transition seems to display mixed aspects of a first-order transition, evidenced by a discontinuity in the coordination number, and a second-order transition, indicated by power-law scalings and diverging lengths. Here we demonstrate that jamming is a first-order transition with quenched disorder in cyclically sheared systems with quasistatic deformations, in two and three dimensions. Based on scaling analyses, we show that fluctuations of the jamming density in finite-sized systems have important consequences on the finite-size effects of various quantities, resulting in a square relationship between disconnected and connected susceptibilities, a key signature of the first-order transition with quenched disorder. This study puts the jamming transition into the category of a broad class of transitions in disordered systems where sample-to-sample fluctuations dominate over thermal fluctuations, suggesting that the nature and behavior of the jamming transition can be understood within the developed theoretical framework of the athermally driven random-field Ising model.
Introduction.
@@
(i) Mechanical marginality, related diverging length scales and power-law scalings. At the jamming transition density (volume fraction) $\varphi_{\rm J}$, the isostatic condition needs to be satisfied for the coordination number $Z$ (average number of contacts per particle): $Z=Z_{\rm iso}=2d$ in frictionless, infinite systems, where $d$ is the dimensionality Maxwell (1870); Alexander (1998). Isostaticity implies that at $\varphi_{\rm J}$ the system is marginally stable, inspiring the search for diverging length scales. According to the “cutting argument”, the isostaticity gives rise to a diverging isostatic length scale at jamming, $l^{*}\sim\Delta Z^{-1}\sim(\varphi-\varphi_{\rm J})^{-1/2}$, below which mechanical stability of the bulk system is affected by the cutting boundaries Wyart (2005); wya (2005); Silbert et al. (2005), where $\Delta Z=Z-Z_{\rm iso}$ is the excess coordination number. The effective medium theory gives a scattering length scale diverging at $\varphi_{\rm J}$, $l_{c}\sim\Delta Z^{-1/2}\sim(\varphi-\varphi_{\rm J})^{-1/4}$, below which continuum elasticity breaks down Wyart (2010); DeGiuli et al. (2014a); Lerner et al. (2014). Other related diverging length scales include the transverse wavelength, $\xi_{T}\sim(\varphi-\varphi_{\rm J})^{-\nu_{T}}$ with $\nu_{T}\approx 0.24$ Silbert et al. (2005), and the longitudinal wavelength, $\xi_{L}\sim(\varphi-\varphi_{\rm J})^{-\nu_{L}}$ with $\nu_{L}\approx 0.48$ Silbert et al. (2005).
@@
(iii) Gardner glass phase, landscape marginality and associated criticality. The marginality at jamming has been established by an independent approach within the framework of replica symmetry breaking Charbonneau et al. (2014); Parisi et al. (2020); Charbonneau et al. (2015a); Berthier et al. (2016, 2019); Li et al. (2021); Urbani et al. (2023). Unjammed hard sphere glasses undergo a Gardner transition where the free energy landscape becomes fractal and marginal, and the caging susceptibility diverges Berthier et al. (2016). The entire Gardner phase, including the jamming limit, is critical. In other words, the caging correlation length, $\xi_{\rm G}\sim\infty$, remains infinite near jamming, when the transition is approached ($\varphi\rightarrow\varphi_{\rm J}$) from below ($\varphi<\varphi_{\rm J}$). The mean-field replica theory predicts the values of exponents in the weak force and small gap distributions, $\alpha=0.41269$ and $\theta=0.42311$, coinciding with the relationship $\theta=1/\alpha-2$ given by the mechanical marginal stability analysis Charbonneau et al. (2014, 2015b); Wang et al. (2022). The theory provides an additional scaling relationship between the cage size $\Delta$ and the entropic pressure $p$, $\Delta\sim p^{-\kappa}$ with $\kappa=1.41574$. The exponents appear to be independent of the dimensionality $d$ for $d\geq d_{\rm u}$ Charbonneau et al. (2014, 2015b); Parisi et al. (2018), where $d_{\rm u}=2$ is the conjectured upper critical dimension of the jamming transition Goodrich et al. (2012).
@@
(v) Discontinuity in the coordination number. The coordination number $Z$, which is considered as an order parameter of the jamming transition Liu and Nagel (2010), jumps discontinuously from $Z=0$ below jamming, to $Z\geq Z_{\rm iso}$ above, under quasi-static compression/decompression Makse et al. (2000); O’hern et al. (2003) or shear Vinutha et al. (2020). In unjammed states, $Z=0$ because particles can push each other apart, leaving no overlapping between them. In jammed states, the contact network spans the entire system and the minimum condition to have such networks is $Z=Z_{\rm iso}$. Thus the discontinuous jump in $Z$ at jamming is essentially related to isostaticity. Apparently, as the signature of a first-order transition, this discontinuity is inconsistent with the viewpoint of a continuous transition described above. Note that not all physical quantities exhibit a discontinuity at jamming: for example, the pressure $P$ and the shear modulus $G$ vanish continuously when the jamming transition is approached from above Makse et al. (2000); O’hern et al. (2003). On the other hand, other quantities, such as the bulk modulus $B$ and the fraction of non-rattlers $f_{\rm NR}$, jump abruptly at jamming, similar to $Z$.
@@
Second, in conventional compression protocols, the generated ensemble depends on the initial conditions  Chaudhuri et al. (2010); Ozawa et al. (2012); Jin and Yoshino (2021) and the basins of attraction Bertrand et al. (2016); Ashwin et al. (2012); Martiniani et al. (2017). Here we instead consider an ensemble prepared by cyclic shear, where the states are sampled by well-controlled dynamics similar to those in thermal systems. Recently, the response of amorphous assemblies of particles to cyclic shear has attracted great interest, due to the presence of a nonequilibrium phase transition, called the reversible-irreversible (RI) transition Pine et al. (2005); Corte et al. (2008); Schreck et al. (2013); Nagasawa et al. (2019); Das et al. (2020). Interestingly, the jamming transition lies in the irreversible regime where particle trajectories are asymptotically diffusive Nagasawa et al. (2019); Das et al. (2020). In this study, we restrict our ensemble within the irreversible phase. The jamming density of this ensemble is equivalent to the minimum jamming density, or the jamming-point (J-point) density that is obtained by rapid quench O’hern et al. (2003).
Third, in our ensemble, we carefully exclude partially crystallized and fragile states with $Z<Z_{\rm iso}$ that are sensitive to mechanical perturbations or protocol parameters, and regard them as unjammed states with $Z=0$. If such states are included, the discontinuity in $Z$ diminishes and the jamming transition looks continuous, similar to the results reported in Ref. Heussinger and Barrat (2009) obtained by uniform shear. Thus we expect the impactibility between the criticality viewpoint in (iv) and our first-order picture originating from finite-rate effects. For finite-time scales, the existence of transient states with $Z<Z_{\rm iso}$ can lead to a continuous jamming transition Olsson and Teitel (2007, 2011); Heussinger and Barrat (2009). However, our results suggest that, after a sufficiently long time, any configurations with $Z<Z_{\rm iso}$ would eventually relax to unjammed states with zero energy and inter-particle contacts. Note that the timescale for the system to attain force balance diverges at the shear jamming transition Vinutha et al. (2020), which means that the real quasistatic limit would correspond to extremely small shear rates in large systems.
@@
The above results imply that the configurational space is effectively explored by the dynamical trajectory of the system in the irreversible phase, encouraging us to consider a statistical ensemble generated by shear dynamics. In this study, we fix $\gamma_{\rm max}={0.7}$ unless otherwise specified, and vary $\varphi$ systematically in the window of $[0.833,0.849]$. At each $\varphi$, in total $t_{\rm max}\times N_{\rm s}$ independent configurations are collected to construct the ensemble. In the SI, we present additional results obtained for $\gamma_{\rm max}={1}$. All configurations are typically isotropic since they are collected at $\gamma=0$.
@@
It is well known now that the jamming density $\varphi_{\rm j}$ is not unique and protocol-dependent Parisi and Zamponi (2010); Chaudhuri et al. (2010); Ozawa et al. (2012); Kumar and Luding (2016); Kawasaki and Miyazaki (2020); Das et al. (2020); Babu et al. (2021); Jin and Yoshino (2021) (In this study, $\varphi_{\rm J}$ denotes the J-point density that is the minimum jamming density, and $\varphi_{\rm j}$ denotes the protocol-dependent jamming density): In the compression protocol, $\varphi_{\rm j}$ depends on the compression rate or the density of the initial equilibrium configuration Chaudhuri et al. (2010); Ozawa et al. (2012); in the cyclic athermal quasistatic compression (CAQC) protocol, $\varphi_{\rm j}$ depends on the volumetric strain amplitude (the maximum density $\varphi_{\rm max}$ to which the system is compressed) and the number $N_{\rm c}$ of cycles Kumar and Luding (2016); Kawasaki and Miyazaki (2020); in the CAQS protocol, $\varphi_{\rm j}$ depends on the shear strain amplitude $\gamma_{\rm max}$ and $N_{\rm c}$ Das et al. (2020); Babu et al. (2021). The continuous range of possible $\varphi_{\rm j}$ is called a jamming-line (J-line), and the minimum jamming density on the J-line is the J-point density $\varphi_{\rm J}$ realized by rapid quenching Babu et al. (2021). The protocol-dependence of $\varphi_{\rm j}$ in the current model has been investigated in Ref. Kawasaki and Miyazaki (2020) using the CAQC protocol, and the maximum jamming density obtained there is $\varphi_{\rm j}\approx 0.8465$ after one over-compression cycle, compared to $\varphi_{\rm J}\approx 0.842-0.843$. In order to examine the effects of non-universal $\varphi_{\rm j}$, in this study we prepare two types of samples: rapidly quenched samples with $\varphi_{\rm j}=\varphi_{\rm J}\approx 0.843$, and mechanically trained samples generated by the CAQC protocol with $\varphi_{\rm j}\approx 0.848>\varphi_{\rm J}$.
@@
Previously, the fragile states were obtained by uniform shear at non-zero strains $\gamma>0$ below $\varphi_{\rm J}$ in experiments Bi et al. (2011) and simulations Kumar and Luding (2016). It was proposed that fragile and jammed states differ in the percolation of the strong force network: in fragile states the percolation occurs anisotropically in one direction only, while in jammed states it occurs isotropically in both directions. In this study, all states are collected during cyclic shear at $\gamma=0$, without systematically introducing anisotropy. Indeed, the fragile states with $43<Z<$ 4 can have percolated force networks of non-rattlers in one or two directions (see Fig. 3f and g). Thus in our case, anisotropy cannot effectively distinguish fragile from jammed states. It should be noted that, due to strong fluctuation effects, the fragile states can be generated either before or after jamming during the shear procedure (see Fig. 2e-f).
@@
Because fragile states are unstable, from now on we count them as unjammed states. More specifically, we replace the original distribution $p(Z)$ with the modified distribution $\tilde{p}(Z)$. For simplicity, we omit the tilde and denote $\tilde{p}(Z)$ by $p(Z)$ below.
Figure 4: Stability tests. Average single-particle (a) potential energy $\overline{e}_{i}$ and (b) net force $\overline{f}_{i}$ as functions of $Z$. Lines in (a) and (b) are guides to the eye. (c) Comparison between $p^{\prime}(Z)$ obtained after the compression-decompression perturbation and $\tilde{p}(Z)$ with fragile states counted as unjammed states, for $N=1000$ at $\varphi=0.844$. (d) Fractions of four states as functions of $f_{\rm th}$. Data are obtained for $N=256$ systems at $\varphi=0.841$
, except for (c).
Figure 5: Distribution of the coordination number. (a) Probability distribution $p(Z)$ at a few different $\varphi$ for $N=1000$. (b) $p(Z)$ at $\varphi=0.843$ for a few different $N$. For better visualization, we do not show the unjammed delta-peak at $Z=0$ (see SI Fig. S1b for full distributions). (c) The fraction of jammed states $F_{\rm J}$ as a function of $\varphi$ for a few different $N$. The intersection of curves gives $\varphi_{\rm J}^{\infty}=0.8432$. (d) The data points of $F_{\rm J}$ with different $N$ collapse as a function of $(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}$. The solid line represents fitting to Eq. (4), with two fitting parameters $u=0.041$ and $\sigma_{\varphi}=0.043$. The average coordination number $\langle Z\rangle$ is plotted as a function of $\varphi$ in (e) and of $(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}$ in (f). The solid line in (f) represents $\langle Z\rangle=F_{\rm J}Z_{\rm J}^{*}$ using $u=0.041$, $\sigma_{\varphi}=0.043$ and $Z_{\rm J}^{*}=4.1$. Symbols in (b-f) have the same meanings.
@@
where $\delta(Z)$ and $p_{\rm J}(Z)$ correspond to the unjammed and jammed peaks respectively. The asymptotic jamming density $\varphi_{\rm J}^{\rm\infty}=0.8432$ for $N\rightarrow\infty$ is determined by the intersection of $F_{\rm J}(\varphi)$ curves with different $N$ (see Fig. 5c). This value is consistent with the J-point density $\varphi_{\rm J}\approx 0.842-0.843$ given by previous studies O’hern et al. (2003); Vagberg et al. (2011), which is unambiguously below the maximum protocol-dependent jamming density $\varphi_{\rm j}\approx 0.8465$ reported in Ref. Kawasaki and Miyazaki (2020) and $\varphi_{\rm j}\approx 0.848$ obtained by CAQC in this study. The difference between $\varphi_{\rm J}^{\rm\infty}=0.8432$ by $\gamma_{\rm max}=0.7$ and $\varphi_{\rm J}^{\rm\infty}=0.8435$ by $\gamma_{\rm max}=1$ (see SI Fig. S2b and S3) is too small to conclude any systematic dependence on $\gamma_{\rm max}$. In Ref. Das et al. (2020), an unjamming pocket is reported in the phase diagram for $\gamma_{\rm max}<0.17$, while for larger $\gamma_{\rm max}$, $\varphi_{\rm J}$ is independent of $\gamma_{\rm max}$, consistent with our observations. These results also suggest that the jamming density obtained by cyclic shear is the lowest jamming density on the J-line of all possible jammed states Parisi and Zamponi (2010); Chaudhuri et al. (2010); Babu et al. (2021).
As shown in Fig. 3d, $Z_{\rm J}^{*}$ is independent of $N$ near $\varphi_{\rm J}$, and weakly depends on $\varphi$. In the following scaling analysis, $Z_{\rm J}^{*}$ is approximated by a constant value $Z_{\rm J}^{*}\approx 4.1$ at $\varphi_{\rm J}^{\rm\infty}=0.8432$. To keep the expressions concise, we ignore the corrections to the scaling functions from the $\varphi$-dependence of $Z_{\rm J}^{*}$.
@@
To understand the origin of $\lambda=1/2$ scaling, recall that the athermal jamming transition is not driven by the free energy difference between the two phases. Below we explain the scaling by the scenario of a first-order transition with quenched disorder. For a finite $N$, due to the presence of disorder in the packing structure, the jamming density $\varphi_{\rm J}^{N}$ should fluctuate around the asymptotic value $\varphi_{\rm J}^{\infty}$. Let us assume a simple Gaussian form of the distribution, $\rho(\varphi_{\rm J}^{N})\sim\exp\left[-\frac{(\delta\hat{\varphi}_{\rm J}+u)^{2}}{2\sigma_{\varphi}^{2}}\right],$ where $\delta\hat{\varphi}_{\rm J}=(\varphi_{\rm J}^{N}-\varphi_{\rm J}^{\infty})N^{1/2}$ follows the standard central limit theorem. Note that $\rho(\varphi_{\rm J}^{N})$ has been explicitly measured in the compression protocol O’Hern et al. (2002); O’hern et al. (2003): the width of $\rho(\varphi_{\rm J}^{N})$ scales as $w\sim N^{-0.55}$ in both 2D and 3D, supporting our assumption. The above assumption also predicts, $\varphi_{\rm J}^{\infty}-\varphi_{\rm J}^{N}\sim N^{-1/2}$, independent of the dimensionality. The jammed states are defined by $\varphi>\varphi_{\rm J}^{N}$, and thus
@@
It seems that the non-Gaussian effect in the distribution $\rho(\varphi_{\rm J}^{N})$, which is neglected in the present analysis, is amplified in the data of $\chi_{\rm dis}/\chi_{\rm con}^{2}$, resulting in slight asymmetry. In SI Figs. S3-S6, we show that the scaling functions Eqs. (4)-(7) work in 3D, and for a different max strain $\gamma_{\max}=1$ in 2D. In 3D, we obtain $\varphi_{\rm J}^{\infty}=0.648$, consistent with the previously reported J-point density $\varphi_{\rm J}\approx 0.648$ of this model Das et al. (2020), which is the minimum jamming density on the J-line. For the same 3D model, the protocol-dependent jamming density is obtained up to $\varphi_{\rm j}=0.6616$ by the thermal annealing method Chaudhuri et al. (2010), and up to $\varphi_{\rm j}=0.661$ by the CAQS mechanical training method Das et al. (2020).

### v4

@@ Abstract
Jamming is an athermal transition between flowing and rigid states in amorphous systems such as granular matter, colloidal suspensions, complex fluids and cells. The jamming transition seems to display mixed aspects of a first-order transition, evidenced by a discontinuity in the coordination number, and a second-order transition, indicated by power-law scalings and diverging lengths. Here we demonstrate that jamming is a first-order transition with quenched disorder in cyclically sheared systems with quasistatic deformations, in two and three dimensions. Based on scaling analyses, we show that fluctuations of the jamming density in finite-sized systems have important consequences on the finite-size effects of various quantities, resulting in a square relationship between disconnected and connected susceptibilities, a key signature of the first-order transition with quenched disorder. This study puts the jamming transition into the category of a broad class of transitions in disordered systems where sample-to-sample fluctuations dominate over thermal fluctuations, suggesting that the nature and behavior of the jamming transition might be better understood within the developed theoretical framework of the athermally driven random-field Ising model.
Introduction
@@
(i) Mechanical marginality, related diverging length scales and power-law scalings. At the jamming transition density (volume fraction) $\varphi_{\rm J}$, the isostatic condition needs to be satisfied for the coordination number $Z$ (average number of contacts per particle): $Z=Z_{\rm iso}=2d$ in frictionless, infinite systems, where $d$ is the dimensionality Maxwell (1870); Alexander (1998). Isostaticity implies that at $\varphi_{\rm J}$ the system is marginally stable, inspiring the search for diverging length scales. According to the “cutting argument”, the isostaticity gives rise to a diverging isostatic length scale at jamming, $l^{*}\sim\Delta Z^{-1}\sim(\varphi-\varphi_{\rm J})^{-1/2}$, below which mechanical stability of the bulk system is affected by the cutting boundaries Wyart (2005); wya (2005); Silbert et al. (2005), where $\Delta Z=Z-Z_{\rm iso}$ is the excess coordination number. The effective medium theory gives a scattering length scale diverging at $\varphi_{\rm J}$, $l_{\rm c}\sim\Delta Z^{-1/2}\sim(\varphi-\varphi_{\rm J})^{-1/4}$, below which continuum elasticity breaks down Wyart (2010); DeGiuli et al. (2014a); Lerner et al. (2014). Other related diverging length scales include the transverse wavelength, $\xi_{\rm T}\sim(\varphi-\varphi_{\rm J})^{-\nu_{\rm T}}$ with $\nu_{\rm T}\approx 0.24$ Silbert et al. (2005), and the longitudinal wavelength, $\xi_{\rm L}\sim(\varphi-\varphi_{\rm J})^{-\nu_{\rm L}}$ with $\nu_{\rm L}\approx 0.48$ Silbert et al. (2005).
@@
(iii) Gardner glass phase, landscape marginality and associated criticality. The marginality at jamming has been established by an independent approach within the framework of replica symmetry breaking Charbonneau et al. (2014); Parisi et al. (2020); Charbonneau et al. (2015a); Berthier et al. (2016, 2019); Li et al. (2021); Urbani et al. (2023). Unjammed hard sphere glasses undergo a Gardner transition where the free energy landscape becomes fractal and marginal, and the caging susceptibility diverges Berthier et al. (2016). The entire Gardner phase, including the jamming limit, is critical. In other words, the caging correlation length, $\xi_{\rm G}\sim\infty$, remains infinite near jamming, when the transition is approached ($\varphi\rightarrow\varphi_{\rm J}$) from below ($\varphi<\varphi_{\rm J}$). The mean-field replica theory predicts the values of exponents in the weak force and small gap distributions, $\alpha=0.41\ldots$ and $\theta=0.42\ldots$, coinciding with the relationship $\theta=1/\alpha-2$ given by the mechanical marginal stability analysis Charbonneau et al. (2014, 2015b); Wang et al. (2022). The theory provides an additional scaling relationship between the cage size $\Delta$ and the entropic pressure $p$, $\Delta\sim p^{-\kappa}$ with $\kappa=1.41\ldots$ The exponents appear to be independent of the dimensionality $d$ for $d\geq d_{\rm u}$ Charbonneau et al. (2014, 2015b); Parisi et al. (2018), where $d_{\rm u}=2$ is the conjectured upper critical dimension of the jamming transition Goodrich et al. (2012).
@@
(v) Discontinuity in the coordination number. The coordination number $Z$, which is considered as an order parameter of the jamming transition Liu and Nagel (2010), jumps discontinuously from $Z=0$ below jamming, to $Z\geq Z_{\rm iso}$ above, under quasi-static compression/decompression Makse et al. (2000); O’Hern et al. (2003) or shear Vinutha et al. (2020). In unjammed states, $Z=0$ because particles can push each other apart, leaving no overlapping between them. In jammed states, the contact network spans the entire system and the minimum condition to have such networks is $Z=Z_{\rm iso}$. Thus the discontinuous jump in $Z$ at jamming is essentially related to isostaticity. Apparently, as the signature of a first-order transition, this discontinuity is inconsistent with the viewpoint of a continuous transition described above. Note that not all physical quantities exhibit a discontinuity at jamming: some quantities, such as the pressure $P$ and the shear modulus $G$, vanish continuously when the jamming transition is approached from above Makse et al. (2000); O’Hern et al. (2003) (for example, see Ref. Kumar and Luding (2016)) for the relation between $P$ and $\varphi$). On the other hand, other quantities, such as the bulk modulus $B$ and the fraction of non-rattlers $f_{\rm NR}$, jump abruptly at jamming, similar to $Z$.
@@
Second, in conventional compression protocols, the generated ensemble depends on the initial conditions  Chaudhuri et al. (2010); Ozawa et al. (2012); Jin and Yoshino (2021) and the basins of attraction Bertrand et al. (2016); Ashwin et al. (2012); Martiniani et al. (2017). Here we instead consider an ensemble prepared by cyclic shear, where the states are sampled by well-controlled dynamics similar to those in thermal systems. Recently, the response of amorphous assemblies of particles to cyclic shear has attracted great interest, due to the presence of a nonequilibrium phase transition, called the reversible-irreversible (RI) transition Pine et al. (2005); Corte et al. (2008); Schreck et al. (2013); Nagasawa et al. (2019); Das et al. (2020). Interestingly, the jamming transition lies in the irreversible regime where particle trajectories are asymptotically diffusive Nagasawa et al. (2019); Das et al. (2020). In this study, we restrict our ensemble within the irreversible phase. The jamming density of this ensemble is equivalent to the minimum jamming density, or the jamming-point (J-point) density that is obtained by rapid quench O’Hern et al. (2003). The protocol dependence of the jamming density is not observed in our simulations: the preparation history of the initial condition becomes irrelevant after the first shear cycle, because a steady state is reached before the strain reversal.
Third, in our ensemble, we carefully exclude partially crystallized and fragile states with $Z<Z_{\rm iso}$ that are sensitive to mechanical perturbations or protocol parameters, and regard them as unjammed states with $Z=0$. Any transient and intermediate states generated during shear, which appear before the steady states, are excluded in our analysis. If such states are included, the discontinuity in $Z$ diminishes and the jamming transition looks continuous, similar to the results reported in Ref. Heussinger and Barrat (2009) obtained by uniform shear. Thus we expect the impactibility between the criticality viewpoint in (iv) and our first-order picture originating from finite-rate effects. For finite-time scales, the existence of transient states with $Z<Z_{\rm iso}$ can lead to a continuous jamming transition Olsson and Teitel (2007, 2011); Heussinger and Barrat (2009). However, our results suggest that, after a sufficiently long time, any configurations with $Z<Z_{\rm iso}$ would eventually relax to unjammed states with zero energy and inter-particle contacts. Note that the timescale for the system to attain force balance diverges at the shear jamming transition Vinutha et al. (2020), which means that the real quasistatic limit would correspond to extremely small shear rates in large systems.
@@ Results
The above results imply that the configurational space is effectively explored by the dynamical trajectory of the system in the irreversible phase, encouraging us to consider a statistical ensemble generated by shear dynamics. In this study, we fix $\gamma_{\rm max}={0.7}$ unless otherwise specified, and vary $\varphi$ systematically in the window of $[0.833,0.849]$. At each $\varphi$, in total $t_{\rm max}\times N_{\rm s}$ independent configurations are generated to construct the ensemble. In the SI, we present additional results obtained for $\gamma_{\rm max}={1}$. All configurations are collected at $\gamma=0$, which have weak anisotropy characterized by a non-zero average macroscopic friction coefficient (the ratio between stress and pressure), $\langle\sigma/P\rangle\approx 0.08$.
@@
It is well known now that the jamming density $\varphi_{\rm j}$ is not unique and protocol-dependent Parisi and Zamponi (2010); Chaudhuri et al. (2010); Ozawa et al. (2012); Kumar and Luding (2016); Kawasaki and Miyazaki (2020); Das et al. (2020); Babu et al. (2021); Jin and Yoshino (2021) (In this study, $\varphi_{\rm J}$ denotes the J-point density that is the minimum jamming density, and $\varphi_{\rm j}$ denotes the protocol-dependent jamming density): In the compression protocol, $\varphi_{\rm j}$ depends on the compression rate or the density of the initial equilibrium configuration Chaudhuri et al. (2010); Ozawa et al. (2012); in the cyclic athermal quasistatic compression (CAQC) protocol, $\varphi_{\rm j}$ depends on the volumetric strain amplitude (the maximum density $\varphi_{\rm max}$ to which the system is compressed) and the number $N_{\rm c}$ of cycles Kumar and Luding (2016); Kawasaki and Miyazaki (2020); in the CAQS protocol, $\varphi_{\rm j}$ depends on the shear strain amplitude $\gamma_{\rm max}$ and $N_{\rm c}$ Das et al. (2020); Babu et al. (2021). The continuous range of possible $\varphi_{\rm j}$ is called a jamming-line (J-line), and the minimum jamming density on the J-line is the J-point density $\varphi_{\rm J}$ realized by rapid quenching Babu et al. (2021). The protocol-dependence of $\varphi_{\rm j}$ in the current model has been investigated in Ref. Kawasaki and Miyazaki (2020) using the CAQC protocol, and the maximum jamming density obtained there is $\varphi_{\rm j}\approx 0.8465$ after one over-compression cycle, compared to $\varphi_{\rm J}\approx 0.842-0.843$. In order to examine the effects of non-universal $\varphi_{\rm j}$, in this study we prepare two types of samples: rapidly quenched samples with $\varphi_{\rm j}=\varphi_{\rm J}\approx 0.843$, and mechanically trained samples generated by the CAQC protocol with $\varphi_{\rm j}={0.8479(4)}>\varphi_{\rm J}$.
@@
Previously, the fragile states were obtained by uniform shear at non-zero strains $\gamma>0$ below $\varphi_{\rm J}$ in experiments Bi et al. (2011) and simulations Kumar and Luding (2016). It was proposed that fragile and jammed states differ in the percolation of the strong force network: in fragile states the percolation occurs anisotropically in one direction only, while in jammed states it occurs isotropically in both directions. In this study, all states are collected during cyclic shear at $\gamma=0$, and the fragile states with $43<Z<$ 4 can have percolated force networks of non-rattlers in one or two directions (see Fig. 3f and g). Thus in our case, anisotropy cannot effectively distinguish fragile from jammed states. It should be noted that, due to strong fluctuation effects, the fragile states can be generated either before or after jamming during the shear procedure (see Fig. 2e-f).
@@
Because fragile (and partially crystallized) states with $43<Z<$ 4 are unstable, from now on we count them as unjammed states. More specifically, we replace the original distribution $p(Z)$ with the modified distribution $\tilde{p}(Z)$. For simplicity, we omit the tilde and denote $\tilde{p}(Z)$ by $p(Z)$ below.
Figure 4: Stability tests. Average single-particle (a) potential energy $\overline{e}_{i}$ and (b) net force $\overline{f}_{i}$ as functions of $Z$. Lines in (a) and (b) are guides to the eye. (c) Comparison between $p^{\prime}(Z)$ obtained after the compression-decompression perturbation and $\tilde{p}(Z)$ with fragile states counted as unjammed states, for $N=1000$ at $\varphi=0.844$. (d) Fractions of four states as functions of $f_{\rm th}$. Data are obtained for $N=256$ systems at $\varphi=0.841$, except for (c).
Figure 5: Distribution of the coordination number. (a) Probability distribution $p(Z)$ at a few different $\varphi$ for $N=1000$. (b) $p(Z)$ at $\varphi=0.843$ for a few different $N$. For better visualization, we do not show the unjammed delta-peak at $Z=0$ (see SI Fig. S1b for full distributions). (c) The fraction of jammed states $F_{\rm J}$ as a function of $\varphi$ for a few different $N$. The intersection of curves gives $\varphi_{\rm J}^{\infty}={0.8432(2)}$. (d) The data points of $F_{\rm J}$ with different $N$ collapse as a function of $(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}$. The solid line represents fitting to Eq. (4), with two fitting parameters $u={0.041(1)}$ and $\sigma_{\varphi}={0.043(1)}$. The average coordination number $\langle Z\rangle$ is plotted as a function of $\varphi$ in (e) and of $(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}$ in (f). The solid line in (f) represents $\langle Z\rangle=F_{\rm J}Z_{\rm J}^{*}$ using $u=0.041$, $\sigma_{\varphi}=0.043$ and $Z_{\rm J}^{*}=4.1$. Symbols in (b-f) have the same meanings.
@@
where $\delta(Z)$ and $p_{\rm J}(Z)$ correspond to the unjammed and jammed peaks respectively. The asymptotic jamming density $\varphi_{\rm J}^{\rm\infty}={0.8432(2)}$ for $N\rightarrow\infty$ is determined by the intersection of $F_{\rm J}(\varphi)$ curves with different $N$ (see Fig. 5c). This value is consistent with the J-point density $\varphi_{\rm J}\approx 0.842-0.843$ given by previous studies O’Hern et al. (2003); Vagberg et al. (2011), which is unambiguously below the maximum protocol-dependent jamming density $\varphi_{\rm j}\approx 0.8465$ reported in Ref. Kawasaki and Miyazaki (2020) and $\varphi_{\rm j}\approx 0.848$ obtained by CAQC in this study. The difference between $\varphi_{\rm J}^{\rm\infty}={0.8432(2)}$ by $\gamma_{\rm max}=0.7$ and $\varphi_{\rm J}^{\rm\infty}={0.8435(1)}$ by $\gamma_{\rm max}=1$ (see SI Fig. S2b and S3-4) is too small to conclude any systematic dependence on $\gamma_{\rm max}$. In Ref. Das et al. (2020), an unjamming pocket is reported in the phase diagram for $\gamma_{\rm max}<0.17$, while for larger $\gamma_{\rm max}$, $\varphi_{\rm J}$ is independent of $\gamma_{\rm max}$, consistent with our observations. These results also suggest that the jamming density obtained by cyclic shear is the lowest jamming density on the J-line of all possible jammed states Parisi and Zamponi (2010); Chaudhuri et al. (2010); Babu et al. (2021). We emphasize that $\gamma_{\rm max}=0.7$ in this study is significantly larger than the yield strain $\gamma_{\rm Y}\approx 0.1$ (see SI Fig. S5a). For a small strain amplitude, $\gamma_{\rm max}=0.05$, we observe a logarithmic growth of $\varphi$ as a function of $t$ under a constant pressure condition Knight et al. (1995); Kumar and Luding (2016), suggesting that the jamming density can be increased by cyclic shear with a small $\gamma_{\rm max}$ Das et al. (2020); Babu et al. (2021); in contrast, for $\gamma_{\rm max}=0.7$, $\varphi$ remains as a constant (see SI Fig. S5b)
As shown in Fig. 3d, $Z_{\rm J}^{*}$ is independent of $N$ near $\varphi_{\rm J}$, and weakly depends on $\varphi$. In the following scaling analysis, $Z_{\rm J}^{*}$ is approximated by a constant value $Z_{\rm J}^{*}{=4.1(1)}$ at $\varphi_{\rm J}^{\rm\infty}=0.8432$. To keep the expressions concise, we ignore the corrections to the scaling functions from the $\varphi$-dependence of $Z_{\rm J}^{*}$.
@@
To understand the origin of $\lambda=1/2$ scaling, recall that the athermal jamming transition is not driven by the free energy difference between the two phases. Below we explain the scaling by the scenario of a first-order transition with quenched disorder. For a finite $N$, due to the presence of disorder in the packing structure, the jamming density $\varphi_{\rm J}^{N}$ should fluctuate around the asymptotic value $\varphi_{\rm J}^{\infty}$. Let us assume a simple Gaussian form of the distribution, $\rho(\varphi_{\rm J}^{N})\sim\exp\left[-\frac{(\delta\hat{\varphi}_{\rm J}+u)^{2}}{2\sigma_{\varphi}^{2}}\right],$ where $\delta\hat{\varphi}_{\rm J}=(\varphi_{\rm J}^{N}-\varphi_{\rm J}^{\infty})N^{1/2}$ follows the standard central limit theorem. Note that $\rho(\varphi_{\rm J}^{N})$ has been explicitly measured in the compression protocol O’Hern et al. (2002, 2003): the width of $\rho(\varphi_{\rm J}^{N})$ scales as $w\sim N^{-0.55}$ in both 2D and 3D, supporting our assumption. The above assumption also predicts, $\varphi_{\rm J}^{\infty}-\varphi_{\rm J}^{N}\sim N^{-1/2}$, independent of the dimensionality. The jammed states are defined by $\varphi>\varphi_{\rm J}^{N}$, and thus
@@
It seems that the non-Gaussian effect in the distribution $\rho(\varphi_{\rm J}^{N})$, which is neglected in the present analysis, is amplified in the data of $\chi_{\rm dis}/\chi_{\rm con}^{2}$, resulting in slight asymmetry. In SI Figs. S3-4 and S6-7, we show that the scaling functions Eqs. (4)-(7) work in 3D, and for a different max strain $\gamma_{\max}=1$ in 2D. In 3D, we obtain $\varphi_{\rm J}^{\infty}={0.6479(2)}$, consistent with the previously reported J-point density $\varphi_{\rm J}\approx 0.648$ of this model Das et al. (2020), which is the minimum jamming density on the J-line. For the same 3D model, the protocol-dependent jamming density is obtained up to $\varphi_{\rm j}=0.6616$ by the thermal annealing method Chaudhuri et al. (2010), and up to $\varphi_{\rm j}=0.661$ by the CAQS mechanical training method Das et al. (2020).